OpenStudy (anonymous):

x(x-2)y'+2y+=0 by using linear differential equation

OpenStudy (anonymous):

divide by x(x-2)

OpenStudy (anonymous):

\[y'+\frac{2}{x^2-2x}y=0\]

OpenStudy (anonymous):

find the integration factor \[\mu(t)\]

OpenStudy (anonymous):

what's the p(x) and Q(x)?

OpenStudy (anonymous):

there is no q(x) since this is an order 1 linear

OpenStudy (anonymous):

\[p(x)=\frac{2}{x^2-2x}\]

OpenStudy (anonymous):

should i use partial fraction ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

don't multiply it like i did

OpenStudy (anonymous):

ok ,

OpenStudy (anonymous):

so the answer must be y(x-2)=cx?

OpenStudy (anonymous):

what did you get from the integration factor

OpenStudy (anonymous):